# Gaussian-Process-based Adaptive Tracking Control with Dynamic Active Learning for Autonomous Ground Vehicles

### v1

@@ IV-B Decoupling
$\underbrace{\begin{bmatrix}\dot{q}\\ \dot{e}_{s}\\ \ddot{e}_{s}\end{bmatrix}}_{\dot{\raisebox{2.0pt}{$\chi$}}_{\mathrm{la}}}=$ $\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&-\frac{C_{\mathrm{f}}+C_{\mathrm{r}}}{mv_{\xi}}\end{bmatrix}}_{A_{\mathrm{la}}(v_{\xi})}\underbrace{\begin{bmatrix}q\\ e_{s}\\ \dot{e}_{s}\end{bmatrix}}_{\raisebox{2.0pt}{$\chi$}_{\mathrm{la}}}+\underbrace{\begin{bmatrix}0\\ 0\\ \frac{C_{\mathrm{f}}}{m}\end{bmatrix}}_{B_{\mathrm{la}}}\delta$ (20)
@@
$+$ $\underbrace{\begin{bmatrix}\!0\!\\ 0\\ \!\frac{l_{\mathrm{r}}C_{\mathrm{r}}-l_{\mathrm{f}}C_{\mathrm{f}}}{m}-1\!\end{bmatrix}}_{B_{c}}\underbrace{c(s)}_{w_{c}}+\underbrace{\begin{bmatrix}0\!\!&\!\!0\!\!\\ 0\!\!&\!\!0\!\!\\ \frac{C_{\mathrm{r}}+C_{\mathrm{f}}}{m}\!\!&\!\!-\frac{l_{\mathrm{f}}^{2}C_{\mathrm{f}}+l_{\mathrm{r}}^{2}C_{\mathrm{r}}}{I_{z}v_{\xi}}\end{bmatrix}}_{B_{2}(v_{\xi})}\underbrace{\begin{bmatrix}\theta_{\mathrm{e}}\\ \dot{\theta}_{\mathrm{e}}\end{bmatrix},}_{w_{2}}$
[…]
@@ IV-C GP-based Model Augmentation
[…]
@@ IV-D Adaptive Control Design
$d_{\mathcal{GP}}$ $=1/B_{\mathrm{lo}}(\delta)\mu_{\mathrm{lo}}(z),$ (22a)
$\delta_{\mathcal{GP}}$ $=B_{\mathrm{la}}^{\dagger}B_{\mathcal{GP}}\mu_{\mathrm{la}}(z).$ (22b)
The adaptivity of (22) comes from the recursive estimation and compensation with the GPs. Furthermore, the posterior variance characterizing the uncertainty of the GP approximation is utilized to systematically collect informative training data and thus improve the sample efficiency of learning [15].
@@
$\delta_{\mathrm{nom}}=$ $K_{\mathrm{la}}(v_{\xi})\raisebox{2.0pt}{$\chi$}_{\mathrm{la}}-\theta_{\mathrm{e}}-B_{\mathrm{la}}^{\dagger}B_{c}w_{c},$ (23a)
$\begin{split}d_{\mathrm{nom}}=&K_{\mathrm{lo}}(\delta)(v_{\xi}-v^{\mathrm{r}})+1/B_{\mathrm{lo}}(\delta)A_{\mathrm{lo}}(\delta)v^{\mathrm{r}}\\ &-1/B_{\mathrm{lo}}(\delta)w_{0},\end{split}$ (23b)
where $K_{\mathrm{la}}(v_{\xi})$ and $K_{\mathrm{lo}}(\delta)$ are parameter-dependent gain matrices and the additional terms are used to achieve reference tracking and to cancel out known disturbances. The feedback matrices are obtained using the nominal subsystems (IV-B) and (20) with the LPV-LQR synthesis proposed in [29]. In the latter sections of this work, we refer to (23) as the nominal part of the controller and (23) and (22) combined as the adaptive part of the controller.
[…]
@@
$\begin{bmatrix}-\mathrm{He}(A(\rho)X\!+\!B(\rho)Y(\rho))\par&(\mathcal{Q}X+\mathcal{R}Y(\rho))^{\top}\\ (\mathcal{Q}X+\mathcal{R}Y(\rho))&I\end{bmatrix}$ (25)
[…]
@@
[…]
@@ V-A Problem Statement
The main objective of the experiment design is to collect data points where the approximation of the GP is less reliable, i.e., where the GP needs the most improvement. This problem has been addressed by active learning methods [14, 15], which use the variance of the posterior distribution of the GPs to find new training points. However, these methods cannot be applied in dynamic scenarios such as ours as the input of the GPs corresponds to state variables which means that the vehicle has to be navigated to a certain state configuration for the evaluation. Therefore systematic trajectory planning is required, taking into account the system dynamics. Therefore, we propose an experiment design that incorporates the system dynamics into active learning and directly synthesizes trajectories to explore regions where the variance of the posterior distribution of the GPs is high. Then, completing the training dataset with new data collected along the resulting trajectory, the GPs can be retrained. In the next sections, we consider our specific scenario, where the inputs of the lateral and longitudinal GPs are the velocity states ($z=[v_{\xi}\;v_{\eta}\;\omega]^{\top}$) of the nonlinear model (12). Given baseline GP models, we utilize their posterior distribution to formulate a numerical optimization problem to obtain trajectories for the refinement of the approximation.
@@ V-B Trajectory Parameterization
The spatial trajectories can be easily parameterized with splines, see Sec. III. However, if all spline parameters are considered free variables, the complexity of solving a planning task increases significantly, and imposing shape constraints is difficult. Therefore, we adopt a modified version of the parameterization in [30]. Let an initial trajectory be denoted as $\mathcal{T}_{0}=\{\psi_{0}(s),v_{0}(s)\}$, where $\psi_{0}(s)$ is a two dimensional arc-length parameterized spline curve and $v_{0}(s)$ is the corresponding speed profile. For $\psi_{0}$, we define $n$ number of nodes $\{s_{i}\}_{i=1}^{n}$ along this initial path which can be chosen based on various strategies, e.g. equidistantly. These are depicted with green dots in Fig. 3. Then, we define new waypoints by moving the node by $w_{i}$ along a line perpendicular to the $\mathcal{T}_{0}$ trajectory, where $w_{i}\in[-w_{\mathrm{b}}/2,w_{\mathrm{b}}/2]$ interval. These waypoints are depicted as red crosses. Then, using 2D spline interpolation on the waypoints, we can obtain the path $\psi_{W}(s)$, which is parameterized by $W=\left[w_{1},\dots,w_{n}\right]$ and the initial trajectory $\psi_{0}$, as depicted in Fig. 3.
A similar approach can be used for the speed profiles. We consider a generic initial speed profile $v_{0}(s)$. Then, we define $n$ nodes along $s$ like in the spatial coordinates, depicted in Fig. 3. By perturbing the velocity profile by $\tilde{v}_{i}\in[v_{\mathrm{min}}-v_{0}(s_{i}),{v}_{\mathrm{max}}-v_{0}(s_{i})],i\in\mathbb{I}_{1}^{n}$ at each node, we can first define profile points, then, by using spline interpolation, obtain a new speed profile $v_{V}(s)$ directly parameterized by $V=[v_{0}(s_{1})+\tilde{v}_{1}\dots v_{0}(s_{n})+\tilde{v}_{n}]$. The main benefit of using this parameterization is that a trajectory can be described by $2n$ number of parameters, which is a significant reduction compared to the full spline parameterization. Furthermore, all parameters have clear physical interpretation, which can be used to guarantee that a trajectory is feasible for the vehicle by altering the bounds $w_{\mathrm{b}}$ and $\tilde{v}_{\mathrm{b}}$.
@@ V-C Objective Function Formulation
To formulate the planning task as a numerical optimization problem, we define the objective function as follows. First, we drive a simulated vehicle model with the previously trained controller along a trajectory candidate. The simulation is performed in discrete time, i.e. $x_{k+1}=f_{\mathrm{sim}}(x_{k})$, where $f_{\mathrm{sim}}$ is the nominal nonlinear vehicle dynamics (15), discretized by 4th order Runge-Kutta algorithm. During the simulation at each time instant $k$, we observe the GP inputs $z_{k}=[0_{3\times 3}\;I_{3\times 3}]x_{k}$ and evaluate the variances of the ${\mathcal{GP}}_{\mathrm{lo}}$ and ${\mathcal{GP}}_{\mathrm{la}}$, given by (5b). These variances are denoted as $\hat{\Sigma}_{\mathrm{lo}}(z_{k},X_{k}),\;\hat{\Sigma}_{\mathrm{la}}(z_{k},X_{k})$ to highlight that the variance at $k$ only depends on the current input $z_{k}$ and the inputs of the training dataset $X_{k}$. After the evaluation, we accumulate the variances into a single variable, defined as
$J_{\mathcal{GP}}=\sum_{k=0}^{N_{\mathrm{sim}}}(\hat{\Sigma}_{\mathrm{lo}}(z_{k},X_{k})+\hat{\Sigma}_{\mathrm{la}}(z_{k},X_{k}))$ (27)
Finally, we extend the datasets of the GPs by adding $z_{k}$ at each timestep, i.e. ${X}_{k}={X}_{k-1}\cup z_{k}$, while ${X}_{0}$ is the original input dataset. This approach allows us to formulate an objective function that (a) enforces the vehicle to explore the regions where the GP uncertainty is high, (b) accounts for already explored regions, as the addition of new data points reduces the predictive variance in the neighboring regions of $z_{k}$.
Note that the simulation is only used during the design as the substitute for the real dynamics. After a trajectory has been obtained, the trajectory will be tracked with the real vehicle for the data collection.

### v2

@@ IV-B Decoupling
$\underbrace{\begin{bmatrix}\dot{q}\\ \dot{e}_{s}\\ \ddot{e}_{s}\end{bmatrix}}_{\dot{\raisebox{2.0pt}{$\chi$}}_{\mathrm{la}}}=$ $\underbrace{\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&-\frac{C_{\mathrm{f}}+C_{\mathrm{r}}}{mv_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}\end{bmatrix}}_{A_{\mathrm{la}}(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}})}\underbrace{\begin{bmatrix}q\\ e_{s}\\ \dot{e}_{s}\end{bmatrix}}_{\raisebox{2.0pt}{$\chi$}_{\mathrm{la}}}+\underbrace{\begin{bmatrix}0\\ 0\\ \frac{C_{\mathrm{f}}}{m}\end{bmatrix}}_{B_{\mathrm{la}}}\delta$ (20)
@@
$+$ $\underbrace{\begin{bmatrix}\!0\!\\ 0\\ \!\frac{l_{\mathrm{r}}C_{\mathrm{r}}-l_{\mathrm{f}}C_{\mathrm{f}}}{m}-1\!\end{bmatrix}}_{B_{\mathrm{{{\color[rgb]{0,0,0}\kappa}}}}}\underbrace{{{\color[rgb]{0,0,0}\kappa}}(s)}_{w_{\mathrm{{{\color[rgb]{0,0,0}\kappa}}}}}+\underbrace{\begin{bmatrix}0\!\!&\!\!0\!\!\\ 0\!\!&\!\!0\!\!\\ \frac{C_{\mathrm{r}}+C_{\mathrm{f}}}{m}\!\!&\!\!-\frac{l_{\mathrm{f}}^{2}C_{\mathrm{f}}+l_{\mathrm{r}}^{2}C_{\mathrm{r}}}{I_{\mathrm{z}}v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}\end{bmatrix}}_{B_{2}(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}})}\underbrace{\begin{bmatrix}\theta_{\mathrm{e}}\\ \dot{\theta}_{\mathrm{e}}\end{bmatrix},}_{w_{2}}$
where the state vector $\raisebox{2.0pt}{$\chi$}_{\mathrm{la}}=[q\;e_{s}\;\dot{e}_{s}]^{\top}$ now only contains the lateral error, its integral and derivative, while $A_{\mathrm{la}}(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}})$ is the parameter varying state-transition matrix with scheduling signal $v_{\xi}$ and $B_{\mathrm{la}}$ is the input matrix. Moreover, $w_{\mathrm{{{\color[rgb]{0,0,0}\kappa}}}}$ is the path curvature and $w_{2}$ is used to lump together the unmodeled path and heading dynamics.
@@ IV-C GP-based Model Augmentation
$\begin{split}\dot{\raisebox{2.0pt}{$\chi$}}_{\mathrm{la}}=&A_{\mathrm{la}}(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}})\raisebox{2.0pt}{$\chi$}_{\mathrm{la}}+B_{\mathrm{la}}\delta+B_{\mathrm{{{\color[rgb]{0,0,0}\kappa}}}}w_{\mathrm{{{\color[rgb]{0,0,0}\kappa}}}}+B_{\mathcal{GP}}\mathcal{GP}_{\mathrm{la}}({{\color[rgb]{0,0,0}\mathpzc{x}}}),\end{split}$ (21a)
$\begin{split}\dot{\raisebox{2.0pt}{$\chi$}}_{\mathrm{lo}}=&A_{\mathrm{lo}}(\delta)\raisebox{2.0pt}{$\chi$}_{\mathrm{lo}}+B_{\mathrm{lo}}(\delta)+w_{0}+\mathcal{GP}_{\mathrm{lo}}({{\color[rgb]{0,0,0}\mathpzc{x}}}),\end{split}$ (21b)
where $\mathcal{GP}_{\mathrm{lo}}({{\color[rgb]{0,0,0}\mathpzc{x}}})$ and $\mathcal{GP}_{\mathrm{la}}({{\color[rgb]{0,0,0}\mathpzc{x}}})$ denote the GPs linked to the lateral and longitudinal subsystems, respectively. Based on the structure of the nominal path following model (15), we can observe that the first three equations only capture kinematic relationships. Therefore, we assume that modeling uncertainties only affect the velocity states $v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$, $v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}$, $\omega$, as proposed in [10]. Furthermore, we can also note that $B_{\mathcal{GP}}=[0\;0\;1]^{\top}$ as uncertain dynamic effects only influence $\dot{e}_{s}$.
Observing the original vehicle model (12), we can also note that any change in the environmental conditions affects the dynamics through the acting wheel forces. As these models depend on the velocity states, we choose these variables to construct the GPs. Therefore, the inputs for both the lateral and the longitudinal GPs are ${{\color[rgb]{0,0,0}\mathpzc{x}}}=[v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\;v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\;\omega]^{\top}$.
@@ IV-D Adaptive Control Design
$d_{\mathcal{GP}}$ $=1/B_{\mathrm{lo}}(\delta)\mu_{\mathrm{lo}}({{\color[rgb]{0,0,0}\mathpzc{x}}}),$ (22a)
$\delta_{\mathcal{GP}}$ $=B_{\mathrm{la}}^{\dagger}B_{\mathcal{GP}}\mu_{\mathrm{la}}({{\color[rgb]{0,0,0}\mathpzc{x}}}).$ (22b)
The adaptivity of (22) comes from the recursive estimation and compensation with the GPs. Furthermore, the posterior variance characterizing the uncertainty of the GP approximation is utilized to systematically collect informative training data, which discussed in Sec. V.
@@
$\delta_{\mathrm{nom}}=$ $K_{\mathrm{la}}(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}})\raisebox{2.0pt}{$\chi$}_{\mathrm{la}}-\theta_{\mathrm{e}}-B_{\mathrm{la}}^{\dagger}B_{{\color[rgb]{0,0,0}\kappa}}w_{{\color[rgb]{0,0,0}\kappa}},$ (23a)
$\begin{split}d_{\mathrm{nom}}=&K_{\mathrm{lo}}(\delta)(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-v^{\mathrm{r}})+1/B_{\mathrm{lo}}(\delta)A_{\mathrm{lo}}(\delta)v^{\mathrm{r}}\\ &-1/B_{\mathrm{lo}}(\delta)w_{0},\end{split}$ (23b)
where $K_{\mathrm{la}}(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}})$ and $K_{\mathrm{lo}}(\delta)$ are parameter-dependent gain matrices and the additional terms are used to achieve reference tracking and to cancel out known disturbances. The feedback matrices are obtained using the nominal subsystems (IV-B) and (20) with the LPV-LQR synthesis proposed in [31]. In the latter sections of this work, we refer to (23) as the nominal part of the controller and (23) and (22) combined as the adaptive part of the controller.
Consider an LPV system in the general form as $\dot{\raisebox{2.0pt}{$\chi$}}=A(\rho)\raisebox{2.0pt}{$\chi$}+B(\rho)u$. The optimal parameter dependent state feedback matrix $K(\rho)$ that minimizes the quadratic cost $J_{\mathrm{LQ}}=\int_{0}^{\infty}\raisebox{2.0pt}{$\chi$}^{\top}\!(t)Q\raisebox{2.0pt}{$\chi$}(t)+u^{\top}\!(t)Ru(t)\;\mathrm{d}t$ with $Q\in\mathbb{R}^{n_{x}\times n_{x}}$ and $R\in\mathbb{R}^{n_{u}\times n_{u}}$ can be obtained by solving the convex optimization problem:
@@
$\begin{bmatrix}-\mathrm{He}(A(\rho)X\!+\!B(\rho)Y(\rho))\par&(\mathcal{Q}X+\mathcal{R}Y(\rho))^{\top}\\ (\mathcal{Q}X+\mathcal{R}Y(\rho))&I\end{bmatrix}$, (25)
where the gain matrices of the quadratic cost are encoded in $\mathcal{Q}=[Q^{\frac{1}{2}}\;0]^{\top}$, $\mathcal{R}=[0\;R^{\frac{1}{2}}]^{\top}$ , and $\mathrm{He}(X)=X^{\top}+X$. Furthermore, $Y(\rho)\in\mathbb{R}^{n_{u}\times n_{x}}$ is parameterized as follows:
@@
and $\mathbb{G}\subset\Gamma$ is a discrete grid of the scheduling region, used to relax the infinite number of LMI constraints. After solving (24), the parameter-dependent feedback matrix is obtained as $K(\rho)=Y(\rho)X^{-1}$. The parameter-dependent state-feedback gains $K_{\mathrm{lo}}$ and $K_{\mathrm{la}}$ for the longitudinal and lateral subsystems have been obtained using the outlined LPV-LQR synthesis with apriori fixed weighting matrices $Q_{\mathrm{lo}}$, $Q_{\mathrm{la}}$, $R_{\mathrm{lo}}$, $R_{\mathrm{la}}$.
@@ V-A Problem Statement
The main objective of the experiment design is to collect data points where the approximation of the GP is less reliable, i.e., where the GP needs the most improvement. This problem has been addressed by active learning methods [14, 15], which use the variance of the posterior distribution of the GPs to find new training points. However, these methods cannot be applied in dynamic scenarios such as ours as the input of the GPs corresponds to state variables which means that the vehicle has to be navigated to a certain state configuration for the evaluation. Thus, systematic trajectory planning is required, taking into account the system dynamics. For this, we propose an experiment design that incorporates the system dynamics into the active learning procedure and directly synthesizes trajectories to explore regions where the variance of the posterior distribution of the GPs is high. Then, completing the training dataset with new data collected along the resulting trajectory, the GPs can be retrained. In the next sections, we consider our specific scenario, where the inputs of the lateral and longitudinal GPs are the velocity states ($z=[v_{\xi}\;v_{\eta}\;\omega]^{\top}$) of the nonlinear model (12). Given baseline GP models, we utilize their posterior distribution to formulate a numerical optimization problem to obtain trajectories for the refinement of the approximation.
@@ V-B Trajectory Parameterization
The spatial trajectories can be easily parameterized with splines, see Sec. III. However, if all spline parameters are considered free variables, the complexity of solving a planning task increases significantly, and imposing shape constraints is difficult. Therefore, we adopt a modified version of the parameterization in [32]. Let an initial trajectory be denoted as $\mathcal{T}_{0}=\{\psi_{0}(s),v_{0}(s)\}$, where $\psi_{0}(s)$ is a two dimensional arc-length parameterized spline curve and $v_{0}(s)$ is the corresponding speed profile. For $\psi_{0}$, we define $n$ number of nodes $\{s_{i}\}_{i=1}^{n}$ along this initial path which can be chosen based on various strategies, e.g. equidistantly. These are depicted with green dots in Fig. 3. Then, we define new waypoints by moving the node by ${\mathpzc{{{\color[rgb]{0,0,0}w}}}}_{i}$ along a line perpendicular to the $\mathcal{T}_{0}$ trajectory, where ${{\color[rgb]{0,0,0}\mathpzc{w}}}_{i}\in[-{{\color[rgb]{0,0,0}\mathpzc{w}}}_{\mathrm{b}}/2,{{\color[rgb]{0,0,0}\mathpzc{w}}}_{\mathrm{b}}/2]$ interval. These waypoints are depicted as red crosses. Then, using 2D spline interpolation on the waypoints, we can obtain the path $\psi_{W}(s)$, which is parameterized by $W=\left[{{\color[rgb]{0,0,0}\mathpzc{w}}}_{1}\dots{\color[rgb]{0,0,0}\mathpzc{w}}_{n}\right]^{\top}$ and the initial trajectory $\psi_{0}$, as depicted in Fig. 3.
A similar approach can be used for the speed profiles. We consider a generic initial speed profile $v_{0}(s)$. Then, we define $n$ nodes along $s$ similarly to the spatial coordinates, depicted in Fig. 3. By perturbing the velocity profile by $\tilde{v}_{i}\in[v_{\mathrm{min}}-v_{0}(s_{i}),{v}_{\mathrm{max}}-v_{0}(s_{i})],\;i\in\mathbb{I}_{1}^{n}$, at each node, we can define profile points. Then, by using spline interpolation, obtain a new speed profile $v_{V}(s)$ directly parameterized by $V=[v_{0}(s_{1})+\tilde{v}_{1}\dots v_{0}(s_{n})+\tilde{v}_{n}]^{\top}$. The main benefit of using this parameterization is that a trajectory can be described by $2n$ number of parameters, which is a significant reduction compared to the full spline parameterization. Furthermore, all parameters have clear physical interpretation, which can be used to guarantee that a trajectory is feasible for the vehicle by altering the bounds ${{\color[rgb]{0,0,0}\mathpzc{w}}}_{\mathrm{b}}$ and $\tilde{v}_{\mathrm{b}}$.
@@ V-C Objective Function Formulation
To formulate the planning task as a numerical optimization problem, we define the objective function as follows. First, we drive a simulated vehicle model with the previously trained controller along a trajectory candidate. The simulation is performed in discrete time, i.e. $x_{k+1}=f_{\mathrm{sim}}(x_{k})$, where $f_{\mathrm{sim}}$ is the nominal nonlinear vehicle dynamics (15), discretized by 4th order Runge-Kutta algorithm. During the simulation, at each time instant $k$, we observe the GP inputs ${{\color[rgb]{0,0,0}\mathpzc{x}}}_{k}=[0_{3\times 3}\;I_{3\times 3}]x_{k}$ and evaluate the variances of the ${\mathcal{GP}}_{\mathrm{lo}}$ and ${\mathcal{GP}}_{\mathrm{la}}$, given by (5b). These variances are denoted as $\hat{\Sigma}_{\mathrm{lo}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{k},X_{k}),\;\hat{\Sigma}_{\mathrm{la}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{k},X_{k})$ to highlight that the variance at $k$ only depends on the current input ${{\color[rgb]{0,0,0}\mathpzc{x}}}_{k}$ and the inputs of the training dataset $X_{k}$. After the evaluation, we accumulate the variances into a single variable, defined as
$J_{\mathcal{GP}}{{\color[rgb]{0,0,0}(V,W)}}=\sum_{k=0}^{N_{\mathrm{sim}}}(\hat{\Sigma}_{\mathrm{lo}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{k},X_{k})+\hat{\Sigma}_{\mathrm{la}}({{\color[rgb]{0,0,0}\mathpzc{x}}}_{k},X_{k})),$ (27)
where the longitudinal and lateral covariances are evaluated by rolling out the controller along a trajectory defined by $V$ and $W$. Finally, we extend the datasets of the GPs by adding ${{\color[rgb]{0,0,0}\mathpzc{x}}}_{k}$ at each timestep, i.e. ${X}_{k}={X}_{k-1}\cup{{\color[rgb]{0,0,0}\mathpzc{x}}}_{k}$, while ${X}_{0}$ is the original input dataset. This approach allows us to formulate an objective function that (a) enforces the vehicle to explore the regions where the GP uncertainty is high, (b) accounts for already explored regions, as the addition of new data points reduces the predictive variance in the neighboring regions of ${{\color[rgb]{0,0,0}\mathpzc{x}}}_{k}$.
Note that the simulation is only used during the design as the substitute for the real dynamics. After a trajectory has been designed, the trajectory will be executed and tracked with the real vehicle for the data collection, which will be used to update the GP.